# SwinSRGAN: Swin Transformer-based Generative Adversarial Network for High-Fidelity Speech Super-Resolution

### v1

@@ Abstract
Speech super-resolution (SR) reconstructs high-frequency content from low-resolution speech signals. Existing systems often suffer from representation mismatch in two-stage mel–vocoder pipelines and from over-smoothing of hallucinated high-band content by CNN-only generators. Diffusion and flow models are computationally expensive, and their robustness across domains and sampling rates remains limited. We propose SwinSRGAN, an end-to-end framework operating on Modified Discrete Cosine Transform (MDCT) magnitudes. It is a Swin Transformer-based U-Net that captures long-range spectro-temporal dependencies with a hybrid adversarial scheme combines time-domain MPD/MSD discriminators with a multi-band MDCT discriminator specialized for the high-frequency band. We employs a sparse-aware regularizer on arcsinh-compressed MDCT to better preserve transient components. The system upsamples inputs at various sampling rates to 48 kHz in a single pass and operates in real time. On standard benchmarks, SwinSRGAN reduces objective error and improves ABX preference scores. In zero-shot tests on HiFi-TTS without fine-tuning, it outperforms NVSR and mdctGAN, demonstrating strong generalization across datasets. Samples can be found online [^1] https://judgeyjj.github.io/.
@@ 1 INTRODUCTION
Early approaches to bandwidth extension predominantly relied on statistical modeling [4, 5] (e.g., codebook-, HMM-, or GMM-based mapping of spectral envelopes). These methods captured coarse spectral trends but struggled with nonstationary and long-range spectro-temporal patterns. With the rise of deep learning, non-generative mapping targeted moderate bandwidths (typically up to 16 kHz): AECNN [6] performed waveform-to-waveform regression, and TFNet [7] jointly modeled time and frequency. Optimizing L1/L2 or STFT objectives recovers formants and low-band harmonics, yet treating SR as deterministic regression over-smooths hallucinated high-band content, limiting perceptual sharpness at 48 kHz. To achieve higher perceptual quality and support higher sampling rates, research shifted to generative modeling. Diffusion- and flow-based approaches (NU-Wave [8], DiffWave [9], WSRGlow [10]) and two-stage pipelines (NVSR [11], AudioSR [12]) leverage strong priors; however, two-stage designs may suffer mel–vocoder representation mismatch and degraded out-of-domain robustness, while diffusion/flow models often incur high computational cost. Building on these foundations, recent advances explore single-step flow matching for efficient generation [13], time-domain SSM-based architectures for end-to-end efficiency [14], and GAN strategies with parallel amplitude–phase prediction [15] or multi-stage sampling control [16]; other lines pursue complex-domain modeling for accurate phase recovery [17] and diffusion-based zero-shot restoration for blind BWE [18].
@@
Motivated by these limitations, vocoder-free SR in an invertible domain avoids the mel–vocoder decoupling. The MDCT-based mdctGAN [19] enables efficient inference on MDCT magnitudes but employs a Pix2Pix-style generator with limited non-local modeling; its arcsinh compression stabilizes training yet suppresses transients, yielding muffled highs and weak band-specific discrimination. In this context, we adopt NVSR and mdctGAN as representative baselines: NVSR typifies strong two-stage mel–vocoder pipelines, and mdctGAN exemplifies vocoder-free, single-stage MDCT-domain modeling; our method leverages their strengths while addressing these limitations.
@@
In this work, we propose SwinSRGAN, a unified end-to-end adversarial framework operating in the MDCT domain. A Swin Transformer-based U-Net generator captures long-range spectro-temporal dependencies while preserving fine-grained details, and a hybrid discriminator integrates time-domain MPD/MSD with a frequency-domain High-Band Multi-Band Discriminator (HB-MBD) to emphasize the reconstructed high-frequency content. The system supports versatile bandwidth extension (from 4–32 kHz to 48 kHz) with real-time inference. Experimental results show consistent improvements over recent speech SR systems on both objective metrics and ABX preference tests. Furthermore, to assess robustness under domain shift, we conduct zero-shot cross-dataset evaluation on HiFi-TTS [20] without fine-tuning; SwinSRGAN consistently achieves lower LSD than NVSR [11] and mdctGAN [19], demonstrating strong generalization beyond the training corpus.
@@
    An end-to-end MDCT-domain SR framework that eliminates mel–vocoder decoupling while remaining vocoder-free;
@@ 2.2 Hybrid Discriminator
As demonstrated in prior work on generative audio models [24, 25], the design of the discriminator framework is critical for generating high-fidelity waveforms. To ensure temporal coherence, we adopt the proven Multi-Period Discriminator (MPD) and Multi-Scale Discriminator (MSD) from HiFi-GAN [24]. The MPD is effective at capturing different periodic patterns in raw audio, while the MSD operates on downsampled versions of the waveform to evaluate structures at various scales. To address this limitation, we adopt a high-frequency–focused High-Band Multi-Band Discriminator (HB-MBD) that operates on MDCT magnitudes. MDCT coefficients are heavy-tailed and become increasingly sparse with frequency; moreover, the upsampling boundary at $f_{lo}=(\text{lr\_sr}/\text{hr\_sr})\,f_{hi}$ defines a region where all coefficients must be hallucinated. To avoid adversarial signal dilution across the intact low band, we place parallel PatchGAN heads only on the added band $f_{lo},f_{hi})$, together with a single full-band head for global shape. Bands are split uniformly in Hz (i.e., equal MDCT-bin width) with a minimum number of bins per head to respect the local stationarity of KBD-windowed MDCT tiles. This design concentrates adversarial pressure where it matters, prevents trading quality between sub-bands, and yields crisper harmonics and fricative energy without destabilizing the low band. where lr_sr and hr_sr denote the input and target sampling rates, respectively; $f_{hi}$ is the Nyquist frequency at hr_sr; and $f_{lo}$ is the corresponding upsampling boundary.
@@ 5 CONCLUSIONS
We presented SwinSRGAN, an end-to-end MDCT-domain SR framework with a Swin Transformer U-Net, a time–frequency hybrid discriminator (MPD/MSD + HB-MBD), and a sparse-aware loss. On VCTK, SwinSRGAN attains the best LSD (0.79) with fewer parameters than most baselines, and ABX tests show listener preference over mdctGAN across input rates. These results verify that modeling long-range spectro-temporal dependencies and focusing adversarial pressure on the added high band improve both detail and fidelity. On a single NVIDIA A100 40 GB GPU, 1 s of wall-clock time produces 5.3 s of 48 kHz audio using our model.

### v2

@@ Abstract
Speech super-resolution (SR) reconstructs high-frequency content from low-resolution speech signals. Existing systems often suffer from representation mismatch in two-stage mel-vocoder pipelines and from over-smoothing of hallucinated high-band content by CNN-only generators. Diffusion and flow models are computationally expensive, and their robustness across domains and sampling rates remains limited. We propose SwinSRGAN, an end-to-end framework operating on Modified Discrete Cosine Transform (MDCT) magnitudes. It is a Swin Transformer-based U-Net that captures long-range spectro-temporal dependencies with a hybrid adversarial scheme combines time-domain MPD/MSD discriminators with a multi-band MDCT discriminator specialized for the high-frequency band. We employs a sparse-aware regularizer on arcsinh-compressed MDCT to better preserve transient components. The system upsamples inputs at various sampling rates to 48 kHz in a single pass and operates in real time. On standard benchmarks, SwinSRGAN reduces objective error and improves ABX preference scores. In zero-shot tests on HiFi-TTS without fine-tuning, it outperforms NVSR and mdctGAN, demonstrating strong generalization across datasets. Samples can be found online [^1] https://judgeyjj.github.io/.
@@ 1 INTRODUCTION
Early approaches to bandwidth extension predominantly relied on statistical modeling [4, 5] (e.g., codebook-, HMM-, or GMM-based mapping of spectral envelopes). These methods captured coarse spectral trends but struggled with nonstationary and long-range spectro-temporal patterns. With the rise of deep learning, non-generative mapping targeted moderate bandwidths (typically up to 16 kHz): AECNN [6] performed waveform-to-waveform regression, and TFNet [7] jointly modeled time and frequency. Optimizing L1/L2 or STFT objectives recovers formants and low-band harmonics, yet treating SR as deterministic regression over-smooths hallucinated high-band content, limiting perceptual sharpness at 48 kHz. To achieve higher perceptual quality and support higher sampling rates, research shifted to generative modeling. Diffusion-based and flow-based approaches (NU-Wave [8], DiffWave [9], WSRGlow [10]) and two-stage pipelines (NVSR [11], AudioSR [12]) leverage strong priors; however, two-stage designs may suffer mel-vocoder representation mismatch and degraded out-of-domain robustness, while diffusion/flow models often incur high computational cost. Building on these foundations, recent advances explore single-step flow matching for efficient generation [13], time-domain SSM-based architectures for end-to-end efficiency [14], and GAN strategies with parallel amplitude-phase prediction [15] or multi-stage sampling control [16]; other lines pursue complex-domain modeling for accurate phase recovery [17] and diffusion-based zero-shot restoration for blind BWE [18].
@@
Motivated by these limitations, vocoder-free SR in an invertible domain avoids the mel-vocoder decoupling. The MDCT-based mdctGAN [19] enables efficient inference on MDCT magnitudes but employs a Pix2Pix-style generator with limited non-local modeling; its arcsinh compression stabilizes training yet suppresses transients, yielding muffled highs and weak band-specific discrimination. In this context, we adopt NVSR and mdctGAN as representative baselines: NVSR typifies strong two-stage mel-vocoder pipelines, and mdctGAN exemplifies vocoder-free, single-stage MDCT-domain modeling; our method leverages their strengths while addressing these limitations.
@@
In this work, we propose SwinSRGAN, a unified end-to-end adversarial framework operating in the MDCT domain. A Swin Transformer-based U-Net generator captures long-range spectro-temporal dependencies while preserving fine-grained details, and a hybrid discriminator integrates time-domain MPD/MSD with a frequency-domain High-Band Multi-Band Discriminator (HB-MBD) to emphasize the reconstructed high-frequency content. The system supports versatile bandwidth extension (from 4-32 kHz to 48 kHz) with real-time inference. Experimental results show consistent improvements over recent speech SR systems on both objective metrics and ABX preference tests. Furthermore, to assess robustness under domain shift, we conduct zero-shot cross-dataset evaluation on HiFi-TTS [20] without fine-tuning; SwinSRGAN consistently achieves lower LSD than NVSR [11] and mdctGAN [19], demonstrating strong generalization beyond the training corpus.
@@
    An end-to-end MDCT-domain SR framework that eliminates mel-vocoder decoupling while remaining vocoder-free;
@@ 2.2 Hybrid Discriminator
As demonstrated in prior work on generative audio models [24, 25], the design of the discriminator framework is critical for generating high-fidelity waveforms. To ensure temporal coherence, we adopt the proven Multi-Period Discriminator (MPD) and Multi-Scale Discriminator (MSD) from HiFi-GAN [24]. The MPD is effective at capturing different periodic patterns in raw audio, while the MSD operates on downsampled versions of the waveform to evaluate structures at various scales. To address this limitation, we adopt a high-frequency-focused High-Band Multi-Band Discriminator (HB-MBD) that operates on MDCT magnitudes. MDCT coefficients are heavy-tailed and become increasingly sparse with frequency; moreover, the upsampling boundary at $f_{lo}=(\text{lr\_sr}/\text{hr\_sr})\,f_{hi}$ defines a region where all coefficients must be hallucinated. To avoid adversarial signal dilution across the intact low band, we place parallel PatchGAN heads only on the added band $f_{lo},f_{hi})$, together with a single full-band head for global shape. Bands are split uniformly in Hz (i.e., equal MDCT-bin width) with a minimum number of bins per head to respect the local stationarity of KBD-windowed MDCT tiles. This design concentrates adversarial pressure where it matters, prevents trading quality between sub-bands, and yields crisper harmonics and fricative energy without destabilizing the low band. where lr_sr and hr_sr denote the input and target sampling rates, respectively; $f_{hi}$ is the Nyquist frequency at hr_sr; and $f_{lo}$ is the corresponding upsampling boundary.
@@ 5 CONCLUSIONS
We presented SwinSRGAN, an end-to-end MDCT-domain SR framework with a Swin Transformer U-Net, a time-frequency hybrid discriminator (MPD/MSD + HB-MBD), and a sparse-aware loss. On VCTK, SwinSRGAN attains the best LSD (0.79) with fewer parameters than most baselines, and ABX tests show listener preference over mdctGAN across input rates. These results verify that modeling long-range spectro-temporal dependencies and focusing adversarial pressure on the added high band improve both detail and fidelity. On a single NVIDIA A100 40 GB GPU, 1 s of wall-clock time produces 5.3 s of 48 kHz audio using our model.